Please help me solve this equation: 3x + 1/3 = 7/5
(or three x plus one third equals seven fifths)
thanks.

Answer by rfer(16322)

3x+1/3=7/5
45/15x+5/15=21/15
45/15x=16/15
x=(16/15)/(45/15)
x=16/15*15/45
x=240/675
x=16/45